OpenStudy (fanduekisses):

How to solve these types of ACT Math Questions? I need help understanding them ( I'm testing tomorrow)

OpenStudy (acxbox22):

are there answer choices?

OpenStudy (acxbox22):

for the first one just multiply 3 x 5 x 6 x 2 to get the total number of outfits

OpenStudy (acxbox22):

number 2 is easy too it gives us the total length so you have to add the ratios 2:5 for the total relationship 2+5=7 21/7=3 so just mutiply 3 by 2 since it is asking for the smaller part so the small is 6 in and the larger part is 15 inches

OpenStudy (phi):

when given a ratio, add the two parts, 2:5 to get the "total" 7 divide that into the actual total 21, to get 3 then multiply 3 by 2 and 5 to get 6:15

OpenStudy (fanduekisses):

what about If x < 0 and (x+3)(x+1)=8, then x^2= A. -5 B. 2 C. 4 D. 5 E. 25

Parth (parthkohli):

Why did you delete your replies...?

OpenStudy (fanduekisses):

because those were already answered and I didn't want people to get confused

OpenStudy (fanduekisses):

I need help with the last one, it seems easy but I don't remember most of pre-algebra to be honest lol